OpenStudy (anonymous):

simplify ([25-x^2]/4x)/([5-x]/x)

OpenStudy (campbell_st):

dividing by a fraction, find the reciprocal of the denominator and multiply \[\frac{25 - x^2}{4x} \times \frac{x}{5-x}\] next factor the quadratic and then take out any common factors \[\frac{(5 - x)(5 + x)}{4x}\times \frac{x}{(5 -x)}\] and the solution becomes obvious.